Eighty pounds of a 54% copper alloy is mixed with 200 lb of a 22% copper
alloy. Find the percent concentration of the resulting mixture. Round to
the nearest tenth of a percent.
1
Expert's answer
2011-02-03T05:00:10-0500
Let's find the mass of copper alloy in the resulting mixture:
80*0.54 + 200*0.22 = 87.2 lb
The total mass of the mixture:
200 + 80 = 280 lb
The concentration of the copper alloy would be
100%*87.2/280 = 31.1 %